has Kc = 4.06 at 500 Celcius . if 0.100 mol of CO and 0.100 mol of H2O are placed in a 1.00-liter reaction vessel at this temperature, what is the concentration of the reactants and products when the system reaches equilibrium?
CO + H2O = CO2 + H2
1
Expert's answer
2020-12-01T09:33:11-0500

Kc= [CO2] x [H2]/[H2O] x [CO]

[CO2]=[H2]= x [H2O] = [CO]=y

C= 0.1 mol/1 L = 0.1 mol/L=0.1M.

Kc= x^2/y^2

y= √(x^2/Kc) =√(0.1^2/4.06)= 0.05 M.

[CO2]=[H2]= 0.05 M

[H2O] = [CO]=0.05M.